Suppose that the economy is characterized by the following behavioral equations, in which all macroeconomic aggregate are measured in billions of Namibian dollars, N$:

C = 160 + 0.6Yd

I = 150

G = 150

T = 100

Solve for

  1. Equilibrium GDP (Y)
  2. Disposable income ( Yd )
  3. Consumption spending ( C )
  4. Multiplier for government expenditure and interpret it.
Answer #1

Answer

a.
AD= C+I+G
AD=Y
Y=160+0.6(Yd)+150+150
Yd=Y-T
Y=160+0.6(Y-T)+150+150
Y=160+0.6(Y-100)+150+150
Y=460+0.6Y-60
0.4Y=520
Y=1300
b.
Yd=Y-T
Yd=1300-100
Yd=1200
c.
C = 160 + 0.6Yd
C=160+0.6(1200)
C=880
d.
m = 1/(1 - c)
c=C/Y
c=880/1300=0.67
m=1/(1-0.67)=3.03

An increase in the government expenditure will create multiple rounds or ipples of stimulus which will lead to an increased effect on GDP.